# Reward-Based Collision-Free Algorithm for Trajectory Planning of Autonomous Robots

### v1

@@ Abstract
This paper introduces a new mission planning algorithm for autonomous robots that enables the reward-based selection of an optimal waypoint sequence from a predefined set. The algorithm computes a feasible trajectory and corresponding control inputs for a robot to navigate between waypoints while avoiding obstacles, maximizing the total reward, and adhering to constraints on state, input and its derivatives, mission time window, and maximum distance. This also solves a generalized prize-collecting traveling salesman problem. The proposed algorithm employs a new genetic algorithm that evolves solution candidates toward the optimal solution based on a fitness function and crossover. During fitness evaluation, a penalty method enforces constraints, and the differential flatness property with clothoid curves efficiently penalizes infeasible trajectories. The Euler spiral method showed promising results for trajectory parameterization compared to minimum snap and jerk polynomials. Due to the discrete exploration space, crossover is performed using a dynamic time-warping-based method and extended convex combination with projection. A mutation step enhances exploration. Results demonstrate the algorithm’s ability to find the optimal waypoint sequence, fulfill constraints, avoid infeasible waypoints, and prioritize high-reward ones. Simulations and experiments with a ground vehicle, quadrotor, and quadruped are presented, complemented by benchmarking and a time-complexity analysis.
@@
This work addresses the common challenge in cyber-physical systems of managing multiple tasks within limited resources, such as available time or energy, where some tasks are more critical or valuable than others, making it essential to allocate and prioritize tasks to maximize total reward within these constraints. We propose an algorithm that optimizes this process in a motion planning scenario, producing trajectories that respect the robot’s capabilities while focusing on higher-priority tasks, potentially bypassing less relevant ones. Our solution is an evolutionary-inspired algorithm that integrates with traditional motion planning methods to select and prioritize tasks based on their importance and resource limitations. This approach optimizes robot trajectories, ensuring they remain within safe limits while addressing top-priority objectives. The algorithm has been tested on various robotic platforms, including a ground vehicle, a robotic dog, and a quadrotor, showcasing its adaptability across systems with different limitations. The practical contributions of this approach lie in its adaptability to several scenarios, where the algorithm can handle different robot platforms, a variety of limitations in term of time, distances, energy, or any combination of these. Currently, the method assumes static reward values and constraints. Future extensions may include real-time changes in the scenario and extends the method for a broader class of robotic platforms, perhaps including direct cooperation between robots and humans.
@@ I Introduction
In recent decades, research in robotics has increasingly focused on enhancing autonomy capabilities, such as for multi-robot missions [1, 2] and human-robot high-level cooperation [3, 4, 5]. In this context, achieving autonomous operations requires motion planning strategies combined with feasible control inputs to guide the robot along the planned trajectory. This process must account for obstacle avoidance and adhere to dynamic and control input constraints, ensuring safe navigation in various environments, and so on.
@@
Existing research has proposed algorithms for trajectory and control planning in static environments, from a starting configuration to a final configuration [6]. Similar algorithms can be employed when there are intermediate tasks as long as they have a predefined order since the problem remains a classical motion planning and control scenario. But when multiple waypoints need to be reached to complete various tasks without a predetermined order, the robot faces the challenge of solving both the motion planning problem and a combinatorial optimization problem. This situation commonly arises in scenarios such as warehouse automation, where robots pick items from diverse locations without a predefined sequence; agricultural robotics, where drones survey and treat different field sections based on dynamic factors like pest outbreaks or crop health; and autonomous delivery systems, where robots efficiently deliver packages to multiple destinations while navigating complex urban environments.
Furthermore, in many real-world applications, the number of available tasks can be substantial, often exceeding the mission’s allowable time, which is typically constrained by factors such as battery capacity. In these situations, it becomes crucial to bypass certain waypoints while maximizing the overall collected reward, as each waypoint or task carries a specific reward value. Fig. 1 illustrates this scenario, where the robot selectively skips less rewarding waypoints to adhere to the distance constraint. Some scenarios includes: drone delivery systems, where a drone with a low battery may prioritize delivering high-priority packages to critical locations while bypassing lower-priority deliveries, search and rescue operations, where robots may need to prioritize assisting the most critically injured individuals first, bypassing less critical waypoints to maximize mission success. Similarly, autonomous underwater vehicles used for environmental monitoring might prioritize collecting data from high-impact areas, such as regions with known pollution or biodiversity hotspots, bypassing less critical waypoints to ensure the collection of the most valuable data before battery depletion. We approach this scenario as a two-layer problem, with the primary layer involving the sequence of waypoints treated as a combinatorial optimization problem, while the secondary layer concentrates on trajectory planning. Accordingly, we review the literature related to these two areas of research.
@@ I-A Literature Review: Combinatorial Optimization
The mission planning problem addressed in this paper can be seen as a variant of the rooted (fixed start) orienteering problem, also known as the bank robber problem or the generalized Traveling Salesman Problem (TSP), with additional constraints involving a maximum mission time window, an obstacle-free path between nodes, and different rewards assigned to each node. This is closely related to the prize-collecting traveling salesman problem. Furthermore, dynamics are added to the traveler as a constraint, including any state and input and its derivatives. Due to the obstacle-free path requirement, the problem can be seen as a bi-level optimization since it involves trajectory planning between a given sequence.
@@
Specifically, given a set of weighted waypoints distributed along a known static environment, the aim is to find a sequence of waypoints and an obstacle-free trajectory between them that maximizes the collected reward subject to specific maximum time and distance values, accounting for energy limitations and mission requirements. The problem is also subject to robot dynamics and control saturation, meaning that even if there is a path between two waypoints, it may involve a non-reachable trajectory due to input saturation or robot dynamics limitations. While the algorithms in the literature [12, 13, 14] provide valuable contributions, they typically do not address weighted waypoints, bypassing waypoints to meet constraints, or integrating dynamics into the motion planning constraints.
@@
Given the constraints, the optimal solution may not collect rewards from all waypoints, leading to solution candidates of varying dimensions. This variability prevents the use of traditional heuristic algorithms, such as particle swarm optimization, or standard crossover techniques in genetic algorithms, as these operations are not well-defined for elements with unmatched dimensions. To determine the optimal waypoint sequence in a variable-dimension space while satisfying dynamic constraints, a genetic algorithm (GA) is proposed and validated. The overall idea of a GA is the evolution of the candidate solutions using evolutionary-inspired rules. Each candidate is evaluated using a fitness function that assigns higher values to the best individuals in the population. Then, based on these values, the next generation emerge from a crossover operation among the previous generation. An extra mutation step is usually added to increase diversity and exploration. This evolution eventually guides the population to high (or low) regions in the design space, finding the global optimal or close in many cases.
@@
In this work, we incorporate the penalty method into the fitness function to account for constraints. During penalty evaluation, the differential flatness property is leveraged to efficiently map each candidate trajectory to its corresponding states and inputs, thereby assessing constraint fulfillment. Moreover, a two-strategy crossover scheme is employed: a subset of the offspring is generated using a crossover that combines dynamic time warping with an extended convex combination and projection, while the remaining subset is produced via the random subsequence insertion method. A similar warping approach is described in [15]. A preliminary obstacle-free path between any two waypoints is calculated during the fitness evaluation using the classical $A^{*}$ algorithm. From the $A^{*}$ intermediate waypoints, a trajectory parameterization method based on clothoid curves is proposed and compared to the classical snap and jerk polynomial methods, showing smoother shapes and curvature adaptability. Further discussion regarding the trajectory generation is presented in Section III-C. Finally, to enhance exploration capabilities and prevent premature convergence, the mutation step can swap two waypoints within the given candidate sequence. The proposal is tested in simulations and validated using a quadrotor, ground vehicle, and a quadruped. A quadruped differential flatness map is proposed, based on the kinematic control strategy.
@@ II Problem Formulation
Consider an environment $\mathcal{E}\subset\mathbb{R}^{2}$, where the elements of $\mathcal{E}$ represent positions. Within this environment, there are $N$ obstacles denoted as $\mathcal{O}_{i}\subset\mathcal{E}$, where the $i$-th obstacle $\mathcal{O}_{i}$ consists of occupied points. Similarly, the obstacle-free space $\mathcal{F}$ is defined as the set difference between the environment and the union of all obstacles, i.e.,
@@ III-A Fitness Function
Step 3: The trajectory is mapped into the state and input space using the DF property. It returns $\boldsymbol{x}(t),\boldsymbol{u}(t)$, and its derivatives, from an open loop control scheme.
Step 4: The state, input, its derivatives, time, and the length of $p(s)$ are compared against the constraint values. The violation measurements $V_{i}(s)$ are estimated based on how large the violations are. Further details are explained in Section III-A2.
@@ III-A2 Constraints Violation Measurement
As explained in Section III-A, the constraints are addressed by the penalty method (6), where each $i$-th constraint violation is measured by $V_{i}$ and weighted by $\alpha_{i}$. The idea is to increase (penalize) the value of the fitness function for those candidates who do not satisfy a constraint, and the penalty is more significant when the violation is greater or holds a longer time.
@@
Next, the obstacle-free constraint should be guaranteed by the trajectory module. Depending on the dimension of the configuration space and the number of waypoints, any trajectory iteration within the fitness evaluation should be avoided to keep the computational cost reasonable. Consequently, we propose a measurement based on the length of the path that it is not in the obstacle-free space $\mathcal{F}$; denoting the obstacle-free constraint as the third constraint:
@@ III-B2 Crossover
Figure 5: Alignment example between the one-dimensional sequences $s_{1}$, $s_{2}$. The optimal alignment of the two sequences is find by the warping path that follows the minimum distance matrix $D$ described by the colorbar.
@@
Notice that, for each entry, the result of (8) lies on the entire space $\mathbb{R}^{d}$. The optimization variable belongs to the finite set (3), therefore a projection to the set $\mathcal{Q}$ is performed for each waypoint. This projection of each element of $s$ can be solved straightforwardly by finding the $\boldsymbol{q}_{i}$ that minimizes the Euclidean norm to that row. Finally, any duplicate waypoint in $s$ is removed, finding $s_{kb}$. This process is illustrated with an example in Fig. 6, where a child sequence $s_{kb}$ is obtained from the crossover of two parent sequences using the described warping, extended convex combination, and projection onto $\mathcal{Q}$.
@@ III-C1 Jerk/Snap Trajectory Challenges
Although most methods in the literature use polynomial splines that minimize jerk or snap, this approach presents an additional challenge beyond what is discussed in Section I-B, which has not been widely recognized in the literature. Since the parameterization is based on time, a set of independent polynomial functions along each coordinate is created $x=P_{x}(t)$, $y=P_{y}(t)$, and such splines can not be represented in arc length form. This cannot be avoided since most trajectories will not fulfill the unique output requirement of the relation $(x,y)$ to be a function (for a coordinate $x$, there could be different $y$ values). Moreover, the curvature can not be controlled directly since the curvature is a relation between $x$ and $y$, but they are independent time polynomials $x=P_{x}(t)$, $y=P_{y}(t)$. The latter means we do not have control under wavy shapes, including sharp knots, which may be undesirable for different reasons: dynamical feasibility, safety, collision avoidance, and efficiency.
@@ III-C2 Clothoid Curves vs Jerk/Snap Trajectories
Figure 9: Clothoid trajectory comparison against minimum snap and jerk (a), and against minimum snap in a high-curvature scenario (b). The clothoid curve may require a larger turning radius but prevents sharp knots and undesired fast-wavy shapes that can be undesired for control systems.
The distance of a given path is calculated using this clothoid curve. Then, the time parameterization is performed depending on the mission time window constraint. Given time and length, an average constant speed is assumed to allocate the time stamps into the path, finally finding the trajectory. This step is justified since the actuator limits are primarily determined by the distribution of time, and as the time goes to infinite, the trajectory approaches a static state that is feasible for systems like quadrotors [20] and ground vehicles. Since we have a mission time window constraint, we use this value as the least aggressive possible trajectory.
@@ IV-A Simulation Experiments
Then, the map from the outputs and its derivatives to the state variables is defined by
@@
and the map from the flat outputs and its derivatives to the input is defined by
@@
Note that we assume the vehicle is translating. Otherwise, $x_{3}$ and $u_{2}$ are not well defined by the diffeomorphism. Moreover, the angle $x_{3}$ has to be consistent with the sign of $u_{1}$, either forward or backward. Now we find the mapping from the states, input, and its derivatives to the output and its derivatives, defined by
@@
Now we consider the same scenario, where one of the waypoints gets a deliberate increment in its weight up to $w=5$. Both results are shown in Fig. 11a and 11b. As a result of the increment of the lower left waypoint reward, the exploration in the upper left zone is reduced to invest that time in reaching the new highly rewarded waypoint. The actual $t_{f}$ of each case ($a$) and ($b$) is 36.8 s and 38.7 s, close to the mission time window, with rewards of 11 and 15, respectively, with no weights assigned to initial and final fixed points.
@@
For the second case of study, consider a map $M_{B}$, where the start and end positions differ, which is typical in scenarios where the robot’s role is to connect a network of points, akin to a transit system connecting multiple distribution centers. The waypoint weights now have different values representing different importance levels of each location. The input constraints are the same as the scenarios on map $M_{A}$. Finally, a mission time window of $40$ s is set. Now, consider the same scenario, with a change in the time constraint to $30$ s, which can happen for mission requirements or just energy efficiency after noting the higher mission time window was unnecessarily large. The results are shown in Fig. 12a and 12b.

### v2

@@ Abstract
This paper proposes a novel mission planning algorithm for autonomous robots that selects an optimal waypoint sequence from a predefined set to maximize total reward while satisfying obstacle avoidance, state, input, derivative, mission time, and distance constraints. The formulation extends the prize-collecting traveling salesman problem. A tailored genetic algorithm evolves candidate solutions using a fitness function, crossover, and mutation, with constraint enforcement via a penalty method. Differential flatness and clothoid curves are employed to penalize infeasible trajectories efficiently, while the Euler spiral method ensures curvature-continuous trajectories with bounded curvature, enhancing dynamic feasibility and mitigating oscillations typical of minimum-jerk and snap parameterizations. Due to the discrete variable length optimization space, crossover is performed using a dynamic time-warping-based method and extended convex combination with projection. The algorithm’s performance is validated through simulations and experiments with a ground vehicle, quadrotor, and quadruped, supported by benchmarking and time-complexity analysis.
@@
This work addresses the challenge of managing multiple tasks within limited resources in cyber-physical systems, where tasks vary in priority and maximizing total reward is critical. We propose an evolutionary-inspired algorithm that optimizes task selection and prioritization in a motion planning context, generating trajectories that respect robot capabilities while favoring higher-priority tasks and potentially bypassing lower-priority ones to fulfill constraints. The algorithm integrates with traditional motion planning techniques, ensuring safe, resource-aware trajectories. It has been validated on a ground vehicle, a quadruped robot, and a quadrotor drone, demonstrating adaptability across platforms with varying constraints. Practical contributions include the method’s flexibility to accommodate different resource limitations, such as time, distance, and energy. The current formulation assumes static rewards and constraints; future work will explore real-time adaptation and extensions to broader classes of robots, including collaborative scenarios involving human-robot interaction.
@@ I Introduction
In recent decades, research in robotics has increasingly focused on enhancing autonomy capabilities, such as for multi-robot missions [1, 2] and high-level human–robot cooperation [3, 4, 5]. In this context, achieving autonomous operations requires motion planning strategies combined with feasible control inputs to guide the robot along the planned trajectory. This process must account for obstacle avoidance and adhere to dynamic and control input constraints, ensuring safe navigation in various environments, and so on.
@@
Existing research has proposed algorithms for trajectory and control planning in static environments, from a starting configuration to a final configuration [6]. Similar algorithms can be employed when there are intermediate tasks, as long as they have a predefined order, since the problem remains a classical motion planning and control scenario. But when multiple waypoints need to be reached to complete various tasks without a predetermined order, the robot faces the challenge of solving both the motion planning problem and a combinatorial optimization problem. This situation commonly arises in scenarios such as warehouse automation, where robots pick items from diverse locations without a predefined sequence; agricultural robotics, where drones survey and treat different field sections based on dynamic factors like pest outbreaks or crop health; and autonomous delivery systems, where robots efficiently deliver packages to multiple destinations while navigating complex urban environments.
Furthermore, in many real-world applications, the number of available tasks can be substantial, often exceeding the mission’s allowable time, which is typically constrained by factors such as battery capacity. In these situations, it becomes crucial to bypass certain waypoints while maximizing the overall collected reward, as each waypoint or task carries a specific reward value. Fig. 1 illustrates this scenario, where the robot selectively skips less rewarding waypoints to adhere to the distance constraint. Some scenarios includes: drone delivery systems, where a drone with a low battery may prioritize delivering high-priority packages to critical locations while bypassing lower-priority deliveries, search and rescue operations, where robots may need to prioritize assisting the most critically injured individuals first, bypassing less critical waypoints to maximize mission success. Similarly, autonomous underwater vehicles used for environmental monitoring might prioritize collecting data from high-impact areas, such as regions with known pollution or biodiversity hotspots, bypassing less critical waypoints to ensure the collection of the most valuable data before battery depletion. We approach this scenario as a two-layer problem, where the primary layer addresses waypoint sequencing as a combinatorial optimization problem, while the secondary layer concentrates on trajectory planning. Accordingly, we review the literature related to these two areas of research.
@@ I-A Literature Review: Combinatorial Optimization
The mission planning problem addressed in this paper can be seen as a variant of the rooted (fixed start) orienteering problem, also known as the bank robber problem or the generalized Traveling Salesman Problem (TSP), with additional constraints involving a maximum mission time window, an obstacle-free path between nodes, and different rewards assigned to each node. This is closely related to the prize-collecting traveling salesman problem. Furthermore, dynamics are added to the traveler as a constraint, including any state and input and their derivatives. Due to the obstacle-free path requirement, the problem can be seen as a bi-level optimization since it involves trajectory planning between a given sequence.
@@
Specifically, given a set of weighted waypoints distributed along a known static environment, the aim is to find a sequence of waypoints and an obstacle-free trajectory between them that maximizes the collected reward subject to specific maximum time and distance values, accounting for energy limitations and mission requirements. The problem is also subject to robot dynamics and control saturation, meaning that even if there is a path between two waypoints, it may involve a non-reachable trajectory due to input saturation or robot dynamics limitations. While the algorithms in the literature [12, 13, 14, 15, 16] provide valuable contributions, they typically do not address weighted waypoints, bypassing waypoints to meet constraints, or integrating dynamics into the motion planning constraints.
@@
Given the constraints, the optimal solution may not collect rewards from all waypoints, leading to solution candidates of varying dimensions. This variability prevents the use of traditional heuristic algorithms, such as particle swarm optimization, or standard crossover techniques in genetic algorithms, as these operations are not well-defined for elements with unmatched dimensions. To determine the optimal waypoint sequence in a variable-dimension space while satisfying dynamic constraints, a genetic algorithm (GA) is proposed and validated. The overall idea of a GA is the evolution of the candidate solutions using evolutionary-inspired rules. Each candidate is evaluated using a fitness function that assigns lower values to the best individuals in the population. Then, based on these values, the next generation emerge from a crossover operation among the previous generation. An extra mutation step is usually added to increase diversity and exploration. This evolution eventually guides the population to high (or low) regions in the design space, finding the global optimal or close in many cases.
@@
In this work, we incorporate the penalty method into the fitness function to account for constraints. During penalty evaluation, the differential flatness property is leveraged to efficiently map each candidate trajectory to its corresponding states and inputs, thereby assessing constraint fulfillment. Moreover, a two-strategy crossover scheme is employed: a subset of the offspring is generated using a crossover that combines dynamic time warping with an extended convex combination and projection, while the remaining subset is produced via the random subsequence insertion method. A similar warping approach is described in [17]. A preliminary obstacle-free path between any two waypoints is calculated during the fitness evaluation using the classical $A^{*}$ algorithm. From the $A^{*}$ intermediate waypoints, a trajectory parameterization method based on clothoid curves is proposed and compared to the classical snap and jerk polynomial methods, showing smoother shapes and curvature adaptability. Further discussion regarding the trajectory generation is presented in Section III-C. Finally, to enhance exploration capabilities and prevent premature convergence, the mutation step can swap two waypoints within a given candidate sequence. The proposal is tested in simulations and validated using a quadrotor, ground vehicle, and a quadruped. A quadruped differential flatness map is proposed, based on the kinematic control strategy.
@@ II Problem Formulation
Consider an environment $\mathcal{E}\subset\mathbb{R}^{2}$, where the elements of $\mathcal{E}$ represent positions. Within this environment, there are $N$ obstacles denoted as $\mathcal{O}_{i}\subset\mathcal{E}$, where the $i$-th obstacle $\mathcal{O}_{i}$ comprises occupied points. Similarly, the obstacle-free space $\mathcal{F}$ is defined as the set difference between the environment and the union of all obstacles, i.e.,
@@ III-A Fitness Function
Step 3: The trajectory is mapped into the state and input space using the DF property. It returns $\boldsymbol{x}(t),\boldsymbol{u}(t)$, and their derivatives, from an open loop control scheme.
Step 4: The state, input, their derivatives, time, and the length of $p(s)$ are compared against the constraint values. The violation measurements $V_{i}(s)$ are computed based on the magnitude and duration of the violations.. Further details are explained in Section III-A2.
@@ III-A2 Constraints Violation Measurement
As explained in Section III-A, the constraints are addressed by the penalty method (6), where each $i$-th constraint violation is measured by $V_{i}$ and weighted by $\alpha_{i}$. The idea is to increase (penalize) the value of the fitness function for those candidates who do not satisfy a constraint, and the penalty is more significant when the violation is greater or holds for a longer time.
@@
Next, the obstacle-free constraint should be guaranteed by the trajectory module. Depending on the dimension of the configuration space and the number of waypoints, any trajectory iteration within the fitness evaluation should be avoided to keep the computational cost reasonable. Consequently, we propose a measurement based on the length of the path that is not in the obstacle-free space $\mathcal{F}$; denoting the obstacle-free constraint as the third constraint:
@@ III-B2 Crossover
Figure 5: Alignment example between the one-dimensional sequences $s_{1}$, $s_{2}$. The optimal alignment of the two sequences is found by the warping path that follows the minimum distance matrix $D$ described by the color bar.
@@
Notice that, for each entry, the result of (8) lies on the entire space $\mathbb{R}^{d}$. The optimization variable belongs to the finite set (3). Therefore, a projection to the set $\mathcal{Q}$ is performed for each waypoint. This projection of each element of $s$ can be solved straightforwardly by finding the $\boldsymbol{q}_{i}$ that minimizes the Euclidean norm to that row. Finally, any duplicate waypoint in $s$ is removed, finding $s_{kb}$. This process is illustrated with an example in Fig. 6, where a child sequence $s_{kb}$ is obtained from the crossover of two parent sequences using the described warping, extended convex combination, and projection onto $\mathcal{Q}$.
@@ III-C1 Jerk/Snap Trajectory Challenges
Although most methods in the literature use polynomial splines that minimize jerk or snap, this approach presents an additional challenge beyond what is discussed in Section I-B, which has not been widely recognized in the literature. Since the parameterization is based on time, a set of independent polynomial functions along each coordinate is created as $x=P_{x}(t)$, $y=P_{y}(t)$, and such splines can not be represented in arc length form. This cannot be avoided since most trajectories will not fulfill the unique output requirement of the relation $(x,y)$ to be a function (for a coordinate $x$, there could be different $y$ values). Moreover, the curvature can not be controlled directly since the curvature is a relation between $x$ and $y$, but they are independent time polynomials $x=P_{x}(t)$, $y=P_{y}(t)$. The latter means we do not have control over wavy shapes, including sharp knots, which may be undesirable for different reasons: dynamical feasibility, safety, collision avoidance, and efficiency.
@@ III-C2 Clothoid Curves vs Jerk/Snap Trajectories
Figure 9: Clothoid trajectory comparison against minimum snap and jerk (a), and against minimum snap in a high-curvature scenario (b). The clothoid curve may require a larger turning radius, but prevents sharp knots and undesired fast-wavy shapes that can be undesired for control systems.
The distance of a given path is calculated using this clothoid curve. Then, the time parameterization is performed depending on the mission time window constraint. Given time and length, an average constant speed is assumed to allocate the timestamps into the path, finally finding the trajectory. This step is justified since the actuator limits are primarily determined by the distribution of time, and as the time goes to infinity, the trajectory approaches a static state that is feasible for systems like quadrotors [22] and ground vehicles. Since we have a mission time window constraint, we use this value as the least aggressive possible trajectory.
@@ III-D Parameter Tuning
To evaluate the sensitivity of each genetic algorithm parameter with respect to the optimization performance, we employed a Random Forest regression model to estimate the relative importance of each parameter in predicting the final cost value.
For each number of waypoints 20, 35, 50 and 65, we trained a separate Random Forest model using the available GA parameter settings as predictors and the resulting BestCost as the response variable. The sensitivity of each parameter was quantified using the out-of-bag permuted predictor importance scores. This method computes the increase in prediction error when the values of a given predictor are randomly permuted, effectively breaking the association between that variable and the response. A larger increase in error indicates a more influential variable.
We repeated this process for each unique number of waypoints. For every parameter, we computed the mean, minimum, and maximum importance scores across problem sizes. These statistics are summarized in the error bar plot shown in Fig. 11, where each marker represents the mean importance and the vertical bars indicate the range between the minimum and maximum values. The results suggest that elitism and the crossover method are the most influential parameters in determining the algorithm’s performance, followed by the truncation parameter. In contrast, the mutation probability exhibits the lowest overall impact. For the population size $c$, we propose a linear growth with respect to the number of waypoints, following the rule $c=\beta_{p}\kappa$ where $\beta_{p}\in[20,40]$, based on similar randomized performance experiments.
Figure 11: Sensitivity of the best solution obtained by the algorithm to variations in genetic algorithm parameters: truncation $T$, elitism $E$, mutation rate $p_{m}$, and crossover method parameter $C$, evaluated across all waypoint configurations. The analysis indicates that elitism and the crossover method are the most influential parameters in determining performance, followed by the truncation parameter. The mutation rate exhibits the least overall impact.
@@ IV-A Simulation Experiments
Then, the map from the outputs and their derivatives to the state variables is defined by
@@
and the map from the flat outputs and their derivatives to the input is defined by
@@
Note that we assume the vehicle is translating. Otherwise, $x_{3}$ and $u_{2}$ are not well defined by the diffeomorphism. Moreover, the angle $x_{3}$ has to be consistent with the sign of $u_{1}$, either forward or backward. Now we find the mapping from the states, input, and their derivatives to the output and its derivatives, defined by
@@
Now we consider the same scenario, where one of the waypoints gets a deliberate increment in its weight up to $w=5$. Both results are shown in Fig. 12a and 12b. As a result of the increment of the lower left waypoint reward, the exploration in the upper left zone is reduced to invest that time in reaching the new highly rewarded waypoint. The actual $t_{f}$ of each case ($a$) and ($b$) is 36.8 s and 38.7 s, close to the mission time window, with rewards of 11 and 15, respectively, with no weights assigned to initial and final fixed points.
@@
For the second case of study, consider a map $M_{B}$, where the start and end positions differ, which is typical in scenarios where the robot’s role is to connect a network of points, akin to a transit system connecting multiple distribution centers. The waypoint weights now have different values representing different importance levels of each location. The input constraints are the same as the scenarios on map $M_{A}$. Finally, a mission time window of $40$ s is set. Now, consider the same scenario, with a change in the time constraint to $30$ s, which can happen for mission requirements or just energy efficiency after noting the higher mission time window was unnecessarily large. The results are shown in Fig. 13a and 13b.